### v2

# Zero-Shot Large Language Models for Long Clinical Text Summarization with Temporal Reasoning

Abstract
Recent advances in large language models (LLMs) have shown potential in clinical text summarization, but their ability to handle long patient trajectories with multi-modal data spread across time remains underexplored. This study systematically evaluates several state-of-the-art open-source LLMs and their Retrieval Augmented Generation (RAG) variants on long-context clinical summarization. We examine their ability to synthesize structured and unstructured Electronic Health Records (EHR) data while reasoning over temporal coherence, by re-engineering existing tasks, including discharge summarization and diagnosis prediction from two publicly available EHR datasets. Our results indicate that long context window improves input integration, but do not consistently enhance clinical reasoning, and LLMs are still struggling with temporal progression and rare disease prediction. While RAG show improvements in hallucination in some cases, it does not fully address these limitations. Our work fills the gap in long clinical text summarization, establishing a foundation for evaluating LLMs with multi-modal data and temporal reasoning.
Zero-Shot Large Language Models for Long Clinical Text Summarization with Temporal Reasoning
Maya Kruse1, Shiyue Hu1,2, Nicholas Derby1,2, Yifu Wu1, Samantha Stonbraker1 Bingsheng Yao3, Dakuo Wang3, Elizabeth Goldberg1, Yanjun Gao1 1University of Colorado Anschutz Medical Campus 2University of Colorado Boulder 3Northeastern University Correspondence: yanjun.gao@cuanschutz.edu
1 Introduction
Figure 1: An illustration of the longitudinal patient trajectory summarization process from multi-modal EHRs. Key causal relationships among medical observations and interventions are highlighted, leading to the final physician diagnosis.
Electronic Health Records (EHRs) encapsulate a wide range of multi-modal data, such as vital signs, laboratory results, radiology findings, and free text clinical notes Mohsen et al. (2022); Li et al. (2022); Belden et al. (2017). Organized across various timestamps, they reflect the dynamic nature of patient care. In clinical settings, particularly for older patients in intensive care units (ICUs) with multiple encounters, chronic conditions, and complex treatment plans, EHRs become especially lengthy and intricate. Clinicians reviewing these long patient histories face significant cognitive burden, often leading to inaccurate decisions such as diagnostic errors Dymek et al. (2021); Singh et al. (2017). Automated summarization can help improve care continuity and decision-making Dymek et al. (2021); Adams et al. (2021); Gao et al. (2023a); Laxmisan et al. (2012); Pivovarov and Elhadad (2015); Liang et al. (2019), but the volume and complexity of EHRs pose challenges, requiring models that handle diverse and evolving clinical data. As in Figure 1, summarizing longitudinal patient trajectories involves extracting temporally ordered clinical events from multi-modal sources. In critical care, robust temporal reasoning is essential to preserve causal relationships among medical events.
Our work evaluates large language models (LLMs) for summarizing complex, longitudinal EHRs, capturing full patient trajectories rather than isolated clinical snapshots. While LLMs have shown promise in clinical NLP tasks Silcox et al. (2024); Wachter and Brynjolfsson (2024); Garcia et al. (2024); Adams et al. (2024); Gao et al. (2022), existing evaluations primarily focus on short-context settings and task-specific fine-tuning. Clinical summarization benchmarks, such as discharge summarization (“Discharge Me”Xu et al. (2024a)) and diagnosis generation from progress notes (“ProbSum”Gao et al. (2023a)), evaluate LLMs on isolated segments of patient records rather than full hospital trajectories.
While techniques like Retrieval-Augmented Generation (RAG) Xu et al. ; Gao et al. (2023b); Lewis et al. (2020) have been explored for diagnosis prediction and discharge summarization Myers et al. (2024); Lyu et al. (2024), their effectiveness in processing long clinical narratives with rich temporal dependencies remains underexplored.
Two key gaps motivate our work: 1) As foundational LLMs with long-context capabilities (over 4k tokens) are adopted into clinical workflows, it is crucial to evaluate their performance without task-specific training to understand their inherent strengths and limitations in processing complex medical information. (2) Existing task settings focus on single time-point records, failing to reflect real-world longitudinal patient care. Our study addresses these gaps by systematically assessing LLMs on full patient trajectories, providing insights into their ability to manage temporally evolving clinical narratives and their applicability in real-world settings.
In this paper, we focus on two public EHR dataset: Medical Information Mart for Intensive Care (MIMIC-III) Johnson et al. (2020) and EHRShot Wornow et al. (2023). We evaluate five state-of-the-art LLMs and their RAG variants: Mistral-7B-Instruct-v0.1  Jiang et al. (2023), Llama3-8B-Instruct  AI@Meta (2024), Qwen2.5-7B  Yang et al. (2024), DeepSeek-R1-Distill-Qwen-32B  DeepSeek-AI et al. (2025) and Llama2-13B-chat-hf  Touvron et al. (2023). Our work contributes to the LLM for clinical text summarization through the following aspects:
  • We re-engineer existing tasks, including discharge summarization, progress note summarization, and diagnosis classification, into new settings for long-context summarization with temporal reasoning, explicitly designed to align with real-world clinical workflows.
  • The tasks allow us to systematically investigate the effects of different modalities and timestamps by incorporating both unstructured clinical notes and structured tabular data.
  • We compare different methods for handling long documents, including directly processing EHR inputs with LLMs and using RAG approaches.
Given the absence of a benchmark for long clinical text summarization that integrates multi-modal data, temporal reasoning, and real-world clinical workflows, our work establishes a foundation for evaluating LLMs in this setting. We provide insights into their strengths and limitations while guiding future research toward more comprehensive, clinically-aligned, and temporally-aware summarization models.
2 Related Work
Clinical text summarization
Existing work includes discharge summarization Xu et al. (2024a); Lyu et al. (2024), diagnosis summarization Gao et al. (2023a, 2022); Liang et al. (2019), hospital course summarization Adams et al. (2021, 2024). While these tasks inherently involve multi-document summarization by human physicians, NLP formulations typically treat them as single-document summarization or multi-document summarization with fixed timestamps (e.g. at the time when patient is discharged).
LLMs for long clinical document
Directly processing the long document input can lead to LLMs’ “lost-in-the-middle" problem Liu et al. (2024). As a result, RAG has been the mainstream when handling long clinical documents, evidenced by their superior performances in diagnosis prediction, discharge summarization and information extraction Myers et al. (2024); Lyu et al. (2024); Lopez et al. (2025). Due to the task setup, these works have only investigated single modality EHRs with long input length.
Our work addresses this problem by extending RAG-based approaches to multi-modal EHR summarization, incorporating both structured (tabular) and unstructured (clinical notes) data to enhance the completeness and accuracy of patient trajectory summaries. We evaluate how well LLMs and their RAG set up can capture and preserve temporal and causal relationships across the full course of a patient’s hospital stay, bridging the gap between isolated document processing and comprehensive longitudinal understanding.
3 Dataset and Tasks
We use two complementary datasets, MIMIC-III and EHRShot, to evaluate LLMs’ clinical reasoning. They differ in three key aspects: 1) Modalities: MIMIC-III includes both structured data and clinical notes, while EHRShot contains only structured data. 2) Output type: MIMIC-III tasks focus on generating short summaries of patient progress or discharge status, whereas EHRShot involves classification-based diagnosis prediction. 3) Decision time span: MIMIC-III targets immediate ICU-related summaries, while EHRShot predicts diagnoses within a year of discharge, emphasizing long-term forecasting based on prior visits. Despite these differences, both datasets address longitudinal patient information, requiring models to reason over time and synthesize complex clinical histories.
3.1 MIMIC-III
MIMIC-III contains data from over 40,000 patients admitted to the ICU at Beth Israel Deaconess Medical Center between 2001 and 2012. Out of all patients, we decided to focus on older patients (65 years and older) who have been admitted to the ICU, accounting for 43.09% of dataset’s total patient population. After consulting with clinicians, we selected patients with a hospital stay exceeding 72 hours from admission, ensuring sufficient accumulated information for summarization while allowing medical uncertainty to stabilize. This specially selected cohort sets our work apart from prior MIMIC-III studies by focusing on older ICU patients who benefit most from accurate LLM-driven summarization, addressing complex cases where clear summaries are crucial for decision-making. By further restricting our selection to hospital stays exceeding 72 hours, we enable a rigorous evaluation of LLMs’ ability to summarize long documents and reason over progressive clinical events, ultimately supporting better care for a population at high risk of adverse outcomes.
MIMIC-III includes physician and nursing notes, ECG reports, and other clinical documents, along with a large amount of structured (tabular) data. We focus on chart events, lab events, input events, and medications, as they are most relevant to patient care. Chart events capture routine vital signs and other bedside measurements, including ventilator settings and mental status. Lab events consist of laboratory test results, such as white blood cell count and cholesterol levels. Input events track administered fluids, including oral/tube feedings and IV infusions, while medications record all drugs given to the patient.
3.1.1 Discharge Summarization
Discharge summaries are a crucial part of a patient’s hospital care process, providing a comprehensive overview of their hospital stay and key clinical events. It is usually written by the physician after the patient is discharged, and contains three main sections: Diagnosis: a list of diagnoses requiring an understanding of the patient’s clinical progression up to discharge; Brief Hospital Course: clinical summary of treatments, interventions, and significant events during hospitalization; Discharge Instructions: post-discharge guidance, including medication plans, dietary recommendations and follow-up care.
Discharge Summarization has been explored in several studies, including  Xu et al. (2024b) and  Ando et al. (2022), often focusing on specific sections like the Brief Hospital Course or discharge instructions. In this paper, we generate all three sections by extracting and chronologically ordering structured and unstructured data from a hospital admission. We limit input to the last 24 hours to prevent overloading the LLM while prioritizing the most relevant information for discharge, considering the generally long hospital stays ($\geq$3 days). We also test a 48-hour window to capture a broader clinical context.
We design four input settings to evaluate LLMs on multi-modal and temporal reasoning: note only where input only contains clinical notes, tabular only where input only contains tabular data, shuffled tabular where we shuffled the tabular data by their timestamps, combined where we combined both notes and tabular data, sorted by their timestamps.
3.1.2 Assessment and Plan (A&P) Generation
Symbol Definition
$D_{i}$ Progress note for day $i$
$X_{i}$ Input data used to generate the progress note for day $i$
Method Input Data for Day $i$
Baseline $X_{i}=\text{EHR}_{i}$ (No prior progress notes)
Single-Day Context $X_{i}=(\text{EHR}_{i},D_{i-1})$ (Includes previous day’s note)
Multi-Day Context $X_{i}=(\text{EHR}_{i},D_{1},D_{2},...,D_{i-1})$ (Includes all previous notes)
Figure 2: Illustration of the A&P generation workflow (top) and a comparison of input data formulations (bottom).
Daily progress notes are the documents where physicians record diagnoses, treatment, and clinical status, providing key insights into a patient’s condition throughout their hospital stay. A progress note typically consists of four sections: Subjective, Objective, Assessment and Plan Weed et al. (1968); Wright et al. (2014). The Subjective and Objective sections serve as the "evidence" and observation of the patient on that day, with Subjective comprising unstructured free text describing symptoms, status, and treatment, while Objective consists of structured tabular data like lab results and charted values. In contrast, the Assessment and Plan sections capture the physician’s reasoning and clinical hypothesis based on this evidence, with diagnoses and treatment plans listed. To assess LLM abilities in summarizing longitudinal data, the Assessment and Plan sections provide a great testbed, requiring integration of information from multiple days, capturing the evolution of the patient’s condition, and synthesizing key clinical evidence into a concise yet comprehensive summary.
Figure 3: The token length of input data used for A&P generation across days and methods (left), as well as that of input data used for Discharge Summarization across time windows and modalities(right).
The setup of the task is illustrated in Figure 2, which explains the three input methods: the no prior context (Baseline) method uses only the current day’s structured EHR data, the Single-Day Context method adds the previous day’s progress note for limited historical context, and the Multi-Day Context method incorporates all prior progress notes for a more comprehensive patient history. For day $i$, the input includes progress notes up to $i-1$ (depending on the method), but the note for $i$ is excluded and compared to the generated version. This setup reflects the real clinical workflow, where physicians reference prior notes when writing new ones.
How long is the input in MIMIC?
Figure 3 shows the input token lengths for both tasks across different settings. Discharge summarization has an average of 500 tokens from the 24 hour and  1940 tokens from the 48 hour window. On Assessment and Plan (A&P) generation, the average token length increases significantly, due to the need to incorporate prior progress notes and structured data over multiple days. The No Context (Baseline) method starts with a relatively high token count on Day 1 but remains lower overall, averaging around 3,125 tokens per day. The Single-Day Context method, which incorporates the previous day’s progress note, leads to a moderate increase in token length, averaging 5,375 tokens. In contrast, the Multi-Day Context method, which accumulates all prior notes, shows the most significant token growth, reaching nearly 10,000 tokens on Day 4 and averaging 6,875 tokens.
3.2 Diagnosis Prediction from EHRShot
The EHRShot dataset  Wornow et al. (2023) is a longitudinal dataset containing only structured EHR data from 6,739 patients at Stanford Medicine, with over 41 million clinical events. Its structured nature makes it distinct from MIMIC, allowing for a more controlled evaluation of models’ ability to process structured temporal information. Additionally, EHRShot includes a broad range of hospital encounters, making it a more comprehensive resource for modeling long-term patient trajectories.
Although our primary focus is on summarizing longitudinal and temporal EHR data, we chose to evaluate models on the six binary diagnosis prediction tasks available in EHRShot. These tasks involve predicting whether a patient will receive a given diagnosis within one year after discharge. While the final prediction is binary, the decision-making process requires synthesizing longitudinal patient history across multiple clinical events, a critical capability for summarization models designed to handle temporal reasoning in EHR data. The selected diagnoses and their prevalence in the dataset are: Hypertension (28.64%), Hyperlipidemia (31.85%), Pancreatic Cancer (20.49%), Celiac (3.46%), Lupus (4.44%), and Acute MI (Myocardial Infarction) (37.04%). The input token lengths vary widely, with a mean of 1,989 tokens ($\sigma$: 1,270) and a median of 1,851. For distinct measurements per patient, the mean is 74 ($\sigma$:48) and a median of 66. More statistics are presented in Appendix A.2.
First entry: Urea Nitrogen is 26 mg/dL. Hematocrit is 34.20%. Hemoglobin is 12.20 g/dL. 12 minutes later: Glucose is 214 mg/dL. Lactate is 2.40 mmol/L. 3 minutes later: Hemoglobin is 13.00 g/dL. Lactate is 2.40 mmol/L. pO2 is 441 mm Hg. 1 hour later: 100.00 ml of 0.9% Normal Saline is administered. Neosynephrine-k is administered. 49 minutes later: Radiology note: 12:48 PM CHEST (PORTABLE AP) Clip Reason: line placement, r/o PTx Admitting Diagnosis: HEAD BLEED; MEDICAL CONDITION: 68-year-old man s/p MVA significant head trauma, intubated s/p r subclavian triple lumen placement…
Table 1: An example of a patient’s compiled data, including relative timestamp (minutes and hours since previous recorded data), structured data converted into narrative format as well as unstructured note data.
3.3 Representing multi-modal, longitudinal patient data
To support the LLMs’ understanding of the structured data, we transform it into natural language, following previous table-to-text manners Gao et al. (2024); Yu et al. (2023). As a first step, the codes for the various measurements in the data are replaced by their natural language description and embedded in a sentence in the form of [Measurement] is [Value][Unit of Measurement]. Input events and medications use the phrase ‘[Measurement] is administered’. Then, after the temporal ordering, the absolute datetime values are replaced by relative values with regard to the preceding timestamp. Measurements taken at the same time are grouped together under one timestamp. An example of a patient chronology consisting of structured data converted to narrative format with a radiology note is given in Table 1.
It is well known that clinical information is frequently recorded and copy-pasted multiple times, a phenomenon commonly referred to as "note bloat," which leads to redundancy and makes it more challenging to extract relevant insights  Shoolin et al. (2013); Liu et al. (2022). This issue is further exacerbated by the large volume of data, which poses challenges when working with LLMs due to their limited context windows. To address this problem and reduce redundancy, we apply preprocessing steps to filter out duplicate records. If a measurement appears in both lab events and chart events, we retain only one instance to prevent duplication. Additionally, when multiple identical records exist within a short time frame, we always retain the latest entry. This approach applies to both structured data and clinical notes, ensuring that copy-pasted or redundant information do not inflate the dataset unnecessarily. On EHRShot data, we remove redundant words from the constructed patient chronologies, and organize data into six-hour chunks based on their timestamp. Similarly to our MIMIC patient cohort, we also focus on patients aged 65 and above, as well as filter out patients with less than 30 measurements. Other details are included in Appendix A.1.
4 Methods
We test five LLMs with various context length: Llama2-13B-chat-hf Touvron et al. (2023) and Mistral-7B-Instruct-v0.1 Jiang et al. (2023), both with a context window of 4K. Additionally, we evaluate Llama3-8B-Instruct AI@Meta (2024), which supports an 8K context window. For LLMs with larger context windows, we test Qwen2.5-7B Yang et al. (2024) and DeepSeek-R1-Distill-Qwen-32B DeepSeek-AI et al. (2025), both capable of handling 128K token input.
4.1 Approaches
We compare a baseline direct generation approach and a structured RAG approach. The key difference between these approaches lies in how they handle long temporal contexts. Direct generation processes the entire input at once but may suffer from the lost-in-the-middle problem, where relevant information buried in long documents is overlooked. In contrast, RAG segments temporal information into retrievable chunks, which helps manage long sequences but may disrupt temporal dependencies between events.
Direct Generation
The patient’s chronological data, formatted according to the specific task requirements, is provided to the model as a direct input without any additional structural modifications. This serves as a baseline to assess the capability of current state-of-the-art language models in processing and generating clinical summaries from raw sequential data, without the aid of task-specific adaptations or architectural enhancements.
RAG
For this setup, we choose the same selection as models used for direct generation, with the exception of DeepSeek, as it is considerably larger than the other LLMs (32B parameters vs 7-13B) yet fails to outperform these smaller models. This combination of high computational cost and suboptimal performance led us to exclude DeepSeek for RAG. Myers et al. (2024) studied the quality of embedding in medical RAG and found BGE Xiao et al. (2023) yielded the highest performance. Thus, we adopt BGE embedding and apply it to all LLM RAG set ups in our work. Additionally, the queries presented in the paper are slightly modified and used to perform query optimization. Hyperparameter optimization is carried out on the standard RAG hyperparameters chunk size, chunk overlap and top-k retrieved documents.
Experiment settings for Direct Generation and RAG
For all LLMs, we run their 8-bit quantized version. We set the output token length as 1,000, but almost all tasks output is significantly shorter than this limit. For RAG setup, we used Langchain Chase (2022) with Faiss Douze et al. (2024) for semantic retrieval in vector database. We perform hyperparameter tuning with chunk size between [250, 750], top k between [10, 50], chunk overlap between [50,200]. The detailed results of hyperparameter searching are in Appendix A.3. All experiments are run on 2 H100 94GB GPUs.
4.2 Evaluation
On MIMIC, where the output is a summary, we report performance using three metrics that capture string overlap, semantics, and concepts. ROUGE-L  Lin (2004) measures string overlap, while BERTScore Zhang et al. (2019) assesses maximum token pairwise similarity. We use SapBERT as the backend for BERTScore due to its superior performance in biomedical entity representation Liu et al. (2020). We also compute the Concept F-score by mapping text to Concept Unique Identifiers (CUIs) using the Unified Medical Language System (UMLS)Lindberg et al. (1993). Specifically, both ground truth summaries and predicted text are converted into CUIs using QuickUMLS, a UMLS-based concept extractorSoldaini and Goharian (2016). We choose these metrics because of findings from prior work:  Croxford et al. (2024) reported that these metrics achieved higher correlation with physician judgement on clinical text evaluation; Myers et al. (2024) reported CUI F-score using QuickUMLS, identifying CUIs from seven semantic types, such as ‘Disease or Syndrome’ and ‘Sign or Symptom’. We adopt the same CUI F-score evaluation as in Myers et al. (2024).
On EHRShot, given that the diagnosis prediction tasks are essentially binary classification, we report macro-average accuracy and F-scores.
5 Results
The results are organized by task: Discharge Summarization and Assessment and Plan Generation performed on MIMIC-III and Diagnosis Prediction on EHRSHOT. Additionally, prompt optimization was carried out on all tasks, using a small sample set (Appendix A.4).
Setting Model CUI ROUGE-L SapBERT Avg.
Direct Gen Mistral 13.46 ±6.41 16.28 ±12.83 65.23 ±13.38 31.66
Llama3 10.47 ±7.88 11.82 ±14.42 55.35 ±22.48 25.88
Qwen 15.13 ±6.44 15.28 ±5.38 63.98 ±13.50 31.46
DeepSeek 14.67 ±7.12 13.51 ±5.04 63.64 ±13.86 30.61
Llama2 15.79 ±7.38 15.34 ±5.90 63.82 ±11.44 31.65
RAG Mistral 12.15 ±2.39 15.04 ±2.34 65.89 ±5.69 31.03
Llama3 16.43 ±5.46 15.90 ±2.00 64.38 ±4.81 32.27
Qwen 16.90 ±4.36 17.91 ±2.35 65.25 ±4.77 33.35
Llama2 15.36 ±4.50 16.98 ±1.59 67.20 ±6.21 33.18
Table 2: Results on discharge summarization, comparing the direct generation and RAG approaches across all models.
Model Sec. CUI ROUGE-L SapBERT
Llama2 Dx 3.87 ± 5.23 4.01 ± 3.17 52.21 ± 11.59
HC 7.07 ± 4.89 12.26 ± 3.50 61.99 ± 6.12
DI 10.28 ± 7.67 13.02 ± 5.18 63.04 ± 11.68
Mistral Dx 5.46 ± 10.07 3.42 ± 4.34 50.07 ± 10.78
HC 8.64 ± 5.98 12.28 ± 3.13 62.51 ± 5.95
DI 9.09 ± 6.90 12.07 ± 3.90 60.05 ± 9.09
Table 3: Discharge Summarization: Llama2 and Mistral Performance Across Sections: Diagnosis (Dx), Hospital Course (HC), and Discharge Instructions (DI).
Model Metric No Prior Single-Day Multi-Day
Direct Generation
Mistral CUI 21.60 ± 6.47 34.06 ± 12.15 30.55 ± 13.48
ROUGE 17.70 ± 3.79 26.11 ± 11.68 23.32 ± 13.05
SapBERT 68.16 ± 6.91 71.63 ± 10.86 72.41 ± 9.47
Average 35.82 ± 5.72 43.93 ± 11.56 42.09 ± 12
Llama3 CUI 21.83 ± 6.51 45.35 ± 10.45 43.55 ± 12.00
ROUGE 16.60 ± 3.39 33.16 ± 10.68 32.42 ± 12.14
SapBERT 72.01 ± 6.04 80.10 ± 8.93 80.42 ± 8.81
Average 36.81 ± 4.80 52.87 ± 10.02 52.13 ± 10.98
Qwen CUI 21.84 ± 5.68 33.8 ± 6.63 33.96 ± 6.56
ROUGE 15.15 ± 3.53 20.74 ± 4.19 21.6 ± 5.23
SapBERT 72.17 ± 4.51 78.01 ± 4.95 78.41 ± 4.81
Average 36.38 ± 4.57 44.18 ± 5.26 44.65 ± 5.53
DeepSeek CUI 22.58 ± 6.63 34.45 ± 9.26 34.15 ± 10.00
ROUGE 16.16 ± 3.15 21.84 ± 5.65 21.07 ± 5.42
SapBERT 74.97 ± 3.66 78.29 ± 4.61 78.35 ± 5.12
Average 37.90 ± 4.48 44.86 ± 6.51 44.52 ± 9.51
RAG
Mistral CUI 22.65 ± 7.94 35.01 ± 12.40 34.17 ± 10.57
ROUGE 17.15 ± 5.60 29.50 ± 12.46 27.40 ± 10.08
SapBERT 68.57 ± 6.16 74.23 ± 7.79 73.55 ± 8.46
Average 36.12 ± 6.57 46.24 ± 10.88 45.04 ± 9.70
Llama3 CUI 20.33 ± 5.59 32.02 ± 5.93 31.60 ± 8.38
ROUGE 15.48± 3.60 22.09 ± 5.51 21.42 ± 6.81
SapBERT 73.78 ± 4.50 78.80 ± 4.96 78.96 ± 5.42
Average 36.53 ± 4.56 44.3 ± 5.47 43.99 ± 6.87
Qwen CUI 22.07± 6.23 37.3 ± 8.83 37.21± 8.15
ROUGE 15.6 ± 3.6 25.42 ± 8.08 25.36 ± 7.5
SapBERT 72.16 ± 4.63 79.27 ± 5.37 80.28 ± 4.68
Average 37.46 ± 4.82 47.33 ± 7.43 47.45 ± 6.78
Table 4: Performance comparison on A&P generation, aggregated over the patient’s entire hospital stay. For readers interested in a more detailed breakdown, we refer them to the Fig 8, where we provide per-day performance results for a subset of patients with a length of stay of at least 5 days.
Figure 4: Average f1 results across modalities and time windows (on direct generation)
Direct Gen RAG
Mistral Qwen Mistral Qwen
Acute MI Accuracy 52.84 62.72 63.03 65.50
F1 33.45 2.58 11.83 16.87
Celiac Disease Accuracy 62.96 96.05 95.04 95.30
F1 2.60 0.00 0.00 9.52
Hyperlipidemia Accuracy 58.02 68.15 66.09 63.03
F1 30.33 1.53 12.74 7.45
Hypertension Accuracy 64.44 69.38 68.81 68.81
F1 26.53 6.06 8.7 8.7
Lupus Accuracy 66.42 95.06 94.79 94.81
F1 10.53 0.00 0.00 0.00
Pancreatic Cancer Accuracy 73.83 78.52 78.86 79.05
F1 13.11 2.25 0.00 0.00
Table 5: Results on the binary diagnosis prediction task using EHRShot data (macro-average).
Discharge Summarization
Table 2 presents results from comparison between the Direct Generation and RAG approaches. Qwen achieves the highest average score of 33.35 using RAG. In general, most LLMs record a performance increase in their RAG variants, the biggest of which is Llama3’s, with an increase of 6.39 on average metric. Mistral is the exception, with a minor performance decrease of -0.63. DeepSeek, as the largest LLM, reports mediocre performance.
Table 3 presents the section-wise analysis of Llama2 and Mistral on direct generation, the two best LLMs directly handling long input. Diagnosis prediction, which demands the most clinical reasoning, shows the worst performance for both LLMs, highlighting the challenge of reasoning over patient trajectories.
Figure 4 shows the effects of the four modalities (notes + tabular data, notes only, tabular only and shuffled tabular) and two context window sizes (24 and 48 hours) on performance. For the 24 hour window, the pure note modality dominates, performing best on Mistral, Qwen, DeepSeek and Llama2. Llama3 is the outlier and reports the best performance on the combined and shuffled tabular modalities. In the 48h window, trends are less clear, but most models still favor notes.
Temporal order appears unhelpful, as shuffled tabular data performs slightly better than chronological tabular data, likely because the data is near the discharge state, where patients’ clinical conditions stabilize, leading to fewer changes over time.
Assessment and Plan Generation
Table 4 shows that adding prior context (Single-Day or Multi-Day) improves performance across all models, particularly for CUI and ROUGE scores. Llama3 consistently outperforms other models, especially in Single-Day and Multi-Day Contexts, reaching the highest Average scores across both Direct Generation and RAG settings. However, Multi-Day Context does not always yield performance gains over Single-Day, suggesting models struggle with longer patient histories. RAG generally improves performance, particularly for Mistral and Qwen, but its impact varies across models: Llama3 performs better without retrieval.
This task specifically requires clinical reasoning over past captured data (evidence) to summarize patient progression and plan for treatment, yet the relatively low ROUGE and CUI scores indicate that models still struggle with temporal reasoning and integrating historical context effectively.
Figure 5: Correlation of token length and average performance across all tasks and metrics
Figure 6: Expert review scores for Qwen direct generation and RAG on Discharge Summarization
Diagnosis Prediction
We evaluate Mistral and Qwen for this task based on their prior performance. Both models prioritize majority class predictions, leading to inflated accuracy but severely low F1-scores. RAG improves accuracy but does not meaningfully enhance recall or F1-score.
The task remains highly challenging, especially for rare diseases, where models struggle to predict positive cases. Qwen achieves high accuracy but an F-score of zero for celiac disease and lupus, reflecting severe class imbalance. It mostly predicts negatives, occasionally misclassifying positives but failing to identify true cases. Given that 96.54% of celiac and 95.56% of lupus cases are true negatives, its accuracy (96.05% and 95.06%) is barely lower than predicting all negatives.
Overall Performance
Figure 5 illustrate how input length impacts model performance across different LLMs, considering their varying context windows. Qwen (128K context), shows a slight positive correlation, suggesting that longer inputs may improve its performance. In contrast, Mistral and Llama2 (both 4K context), decline as input length increases, likely due to exceeding their optimal processing capacity. Llama3 (8K context), remains stable with minimum impact from the length. Models with shorter contexts struggle with longer inputs, while larger-context models handle them better, though benefits remain inconsistent.
6 Discussion
Another critical aspect to consider is running time and memory usage. All 7B LLMs require approximately 15GB of GPU RAM for direct generation. RAG demands significantly more memory, peaking at 32–33GB, nearly twice that of direct generation. For Qwen, it takes 18-20 minutes to run RAG on one progress note with prior context input, while directly generation takes 10 minutes on average. On EHRShot where input is shorter, diagnosis prediction takes just 3–4 seconds. For other LLMs with smaller context window, they run quicker.
Expert error analysis
A senior board-certified Emergency Department physician was consulted to provide more in-depth analysis of our generated outputs. This assessment was not intended to yield quantitative results but rather to gain insights into how physicians perceive the generated summaries beyond automated metrics and to identify areas for improvement. We focus specifically on discharge summarization and sample 10 pairs of RAG and Direct Generation output from Qwen, as it reached relatively high performance on automated metrics. The physician evaluated the output based on six criteria that have previously been deployed for clinical text evaluation: overall accuracy (factual correctness), hallucination,omission, readability, clinical relevance and specificity Singhal et al. (2023); Xu et al. (2024a); Ben Abacha et al. (2023); Aljamaan et al. (2024); Croxford et al. (2024); Williams et al. (2024). The results of this analysis are given in Figure 6, and the annotation guidelines can be found in Table 15.
Overall, RAG does slightly better, hallucinating less and generating more specific and clinically relevant summaries. However, both approaches still face significant issues: the models tend to focus on less relevant diagnoses and interpret them as the primary reason for admission and oversimplified summaries cause the omission of critical parts of the clinical course. The LLMs also struggle with diagnoses which are stated earlier on but are later discarded or disproved: they fail to realize that these diagnoses are no longer relevant or accurate. This emphasizes the models’ continuing struggle with temporal reasoning.
7 Conclusion
Our work examined LLM’s capabilities on long-context clinical summarization using two publicly available EHR datasets. We found that current LLMs continue to face significant challenges when it comes to accurately summarizing long-context clinical documents. Additionally, these models struggle with temporal reasoning over patient trajectories, making it difficult to track and interpret the sequence of medical events effectively. While RAG has shown some performance gains, the overall results remain substandard, emphasizing the need for further advancements in this area.
Limitations
This study focused on current LLMs’ capabilities on the task of clinical summarization for long context documents including temporal information. We evaluated several different models, including Qwen, Llama 3 and DeepSeek, but we acknowledge that this selection is by no means comprehensive and could have limited our analysis. It also does not contain any closed-source LLMs, as the use of these models with MIMIC and EHRSHOT data is prohibited by Data Use Agreements. Due to the dearth of publicly available EHR datasets, we were only able to include data from two sources: MIMIC-III and EHRSHOT.
Our human evaluation was constrained by real-world clinical scheduling limitations, allowing us to consult only one emergency department (ED) physician. Furthermore, we did not formally validate our proposed survey instrument, though the survey questions were aggregated from established prior work (see Appendix regarding the literatures these questions came from). Our goal is to leverage domain experts to better understand LLM limitations, and we plan to expand human evaluation efforts in future studies to provide a more comprehensive assessment.
Ethical Statement
This study utilizes de-identified patient data from publicly available datasets (MIMIC-III and EHRSHOT), ensuring that no identifiable patient information is used. As a result, our work does not involve human subjects research and poses no risk to patient privacy or confidentiality.
Additionally, our study is purely retrospective and computational, focusing on evaluating LLMs for clinical summarization. The models analyzed do not interact with real-world clinical workflows and are not used for actual medical decision-making. Therefore, there is no potential harm to patients or healthcare providers as a result of this research.
Our goal is to assess and improve LLM capabilities for handling complex clinical data, with the long-term aim of developing safe, explainable, and effective AI tools to support healthcare professionals in the future. However, we acknowledge that applying existing LLMs in clinical workflows carries real risks, including privacy leakage, algorithmic biases, and the potential for incorrect decisions. Addressing these challenges is critical for ensuring the safe and ethical deployment of AI in healthcare.
Appendix A Appendix
A.1 More Preprocessing Details
In addition to the preprocessing steps outlined in section 3.3, we attempt, as far as possible, to only include laboratory measurements and vital sign values that fall outside the normal range. These values are more likely to indicate a problem the patient is facing and are thus more salient. For lab measurements, we only keep those values flagged as ’abnormal’. For chart or vital sign values, there exists a warning flag, which is either 0 or 1. However, some values have an NaN value instead. We exclude values with the flag set to 0, but keep those with 1 or NaN, since it is unknown whether the value corresponding to the NaN type is abnormal or not without a deeper medical analysis.
A.2 EHRSHOT Statistics
Figure 7: Additional statistics on the EHRSHOT dataset. Number of distinct measurements given above, with the token length distribution below.
Figure 7 reports detailed statistics of EHRShot cohort, regarding the input token length and number of distinct clinical measurements. The prompt token lengths exhibit a wide range, with a mean of 1,989 tokens and a standard deviation of 1,270 tokens. The distribution is right-skewed, with a minimum of 178 tokens and a maximum of 5,919 tokens, while the median (50th percentile) is 1,851 tokens.
For distinct measurement types per patient, the data is also highly variable, with a mean of 74 and a standard deviation of 48. The number of distinct measurements ranges from 1 to 351, with a median of 66 and an interquartile range from 39 (25th percentile) to 100 (75th percentile).
Both distributions highlight significant variation in input complexity, reinforcing the need for models to handle long-context dependencies and multi-modal data effectively.
A.3 Hyperparameter searching results
Experiment Top-K Chunk Size Chunk Overlap
Exp 1 10 500 100
Exp 2 20 750 100
Exp 3 50 500 50
Exp 4 20 250 100
Exp 5 50 750 200
Table 6: Hyperparameter configurations for each experiment.
Model Metric Exp 1 Exp 2 Exp 3 Exp 4 Exp 5
Mistral ROUGE-L 6.14 6.18 5.93 6.10 6.37
SapBERT 8.41 8.12 8.37 8.35 8.03
CUI 51.03 50.82 51.12 52.56 51.34
Llama3 ROUGE-L 7.37 7.43 6.95 7.07 6.90
SapBERT 9.23 9.78 9.56 9.19 9.32
CUI 55.45 56.90 56.25 55.89 56.67
Qwen ROUGE-L 6.41 6.97 6.95 6.60 7.11
SapBERT 9.72 9.62 9.68 9.66 10.09
CUI 55.76 56.34 55.42 55.90 56.21
Llama2 ROUGE-L 7.56 7.13 7.59 7.42 6.64
SapBERT 10.22 10.21 9.96 10.16 9.41
CUI 55.67 54.70 54.37 55.80 53.50
Table 7: Hyperparameter Tuning for Discharge Summaries Results (averaged across sections)
Model Metric Exp 1 Exp 2 Exp 3 Exp 4 Exp 5
Mistral ROUGE-L 21.07 17.72 15.74 20.30 15.52
SapBERT 71.42 70.03 68.81 71.45 69.95
CUI 28.05 21.94 20.63 27.92 19.97
Llama3 ROUGE-L 16.73 17.54 16.74 17.11 16.58
SapBERT 71.46 71.66 71.02 71.11 71.60
CUI 22.52 23.88 23.40 23.71 22.50
Qwen ROUGE-L 19.13 21.13 20.39 18.59 20.62
SapBERT 70.29 69.34 70.98 70.22 70.03
CUI 27.28 30.17 28.87 28.04 29.87
Table 8: Hyperparameter Tuning for A&P Generation
In this section, we report our hyperparameter tuning on the RAG setup for all LLMs. Table 6 covers the hyperparameter selection. All tuning are done on Discharge Summarization and A&P generation tasks using a small held-out set (n=5). The best set of parameters are used for official experiments reported in the main text.
A.4 Prompt and query optimization
We optimized the prompts and queries used for RAG systems on all tasks. Both prompts and queries are similar in terms of its instruction, the only difference is that queries for RAG has the “Retrieve" component.
What is the patient’s main diagnosis?
The patient’s primary diagnosis is: (Llama2)
Identify the primary reason for the patient’s hospital admission: (Llama3)
Instruct: Given a search query, retrieve relevant passages that answer the query.
Query: patient’s primary diagnosis. (Mistral)
The patient has been diagnosed with: (Qwen)
Table 9: Queries used for retrieving the primary diagnosis. Highest performing on models bolded with respective model in parentheses.
Summarize the hospital course for this patient in a concise and accurate way.
The patient’s hospital course included the following:
Provide a brief hospital course, including key events and treatments.
(Mistral, Llama3)
Instruct: Given a search query, retrieve relevant passages that answer the query.
Query: Brief hospital course.
What were the key events and outcomes during the patient’s hospital stay?
Table 10: Queries used for retrieving the hospital course. Highest performing on models bolded with respective model in parentheses.
Given the input EHR data, generate discharge instructions for this patient. (All models)
What are the discharge instructions for the patient?
Write a summary of the discharge plan, including medications, follow-up visits, and
patient care instructions.
Instruct: Given a search query, retrieve relevant passages that answer the query.
Query: discharge instructions.
What follow-up care and medications are recommended for the patient after discharge?
Table 11: Queries used for retrieving the discharge instructions. Highest performing on models bolded with respective model in parentheses.
Given the patient EHR data, write the Assessment section of a clinical progress note. The Assessment should include
a brief description of both passive and active diagnoses. Clearly state why the patient is admitted to the hospital
and describe the active problem for the day, along with any relevant comorbidities the patient has.
Provide the Assessment section of the patient’s progress note, including active and passive diagnoses, admission reasons,
the patient’s active problems for the day, and relevant comorbidities.
What are the patient’s active and passive diagnoses? Why was the patient admitted to the hospital?
What are the active medical problems for the day? Include relevant comorbidities.
Retrieve passages that explain the patient’s active and passive diagnoses, reasons for admission, active problems
for the day, and relevant comorbidities. (Mistral, Qwen)
Instruct: Generate a concise Assessment section for the patient’s progress note. Include a summary of active and
passive diagnoses, admission reasons, the patient’s current active problems, and any comorbidities. (Llama3)
Table 12: Queries used for retrieving the Assessment section of a progress note. Highest performing on models bolded with respective model in parentheses.
Given the patient EHR data, write the Plan section of a clinical progress note. The Plan should be organized into multiple
subsections, each corresponding to a specific medical problem. Provide a detailed treatment plan for each problem, outlining
proposed or ongoing interventions, medications, and care strategies.
Generate the Plan section of the patient’s progress note. Provide detailed treatment plans for specific medical problems,
including proposed or ongoing interventions, medications, and care strategies.
What are the proposed treatment plans for the patient’s active medical problems? Include details on interventions,
medication regimens, and care strategies.
Retrieve passages that outline treatment strategies for medical problems, including medications, interventions,
and care strategies.
(Mistral, Qwen)
Instruct: Write the Plan section of the progress note. Organize it into subsections for each medical problem.
Provide detailed plans for treatments, interventions, medications, and care strategies. (Llama3)
Table 13: Queries used for retrieving the Plan section of a progress note. Highest performing on models bolded with respective model in parentheses.
Table 9,10,12,13,11 presents all prompts and queries for MIMIC tasks.
A.5 Impact of context length on consecutive patient data
Figure 8: Impact of prior-day context on F1 score across consecutive days
Figure 8 illustrates the impact of incorporating prior-day context on F1 scores for patient note generation. The x-axis represents consecutive relative days within a patient’s hospital stay, where Day 0 corresponds to the first day of admission and serves as the ground truth (i.e., no evaluation is performed on this day). The y-axis shows the average F1 score across a subset of 10 patients with a length of stay of at least 5 days.
Since Day 0 (Day 1 in the figure) is not evaluated, its F1 score is set to 1.0 for all methods to ensure a clear visual comparison with subsequent days. The results demonstrate that incorporating prior-day context improves performance over time. The Single-Day Context and Multi-Day Context methods achieve substantially higher F1 scores than the No Prior Context method, particularly on Day 2 and Day 3, suggesting that leveraging past information helps generate more accurate and coherent patient notes. However, after Day 3, the performance of the Multi-Day Context method begins to decline, indicating that while longer historical context can be beneficial, it may introduce additional noise or redundant information.
Overall, these findings highlight that considering prior context enhances the accuracy of generated patient notes, with Single-Day Context yielding the highest performance in later days, while Multi-Day Context shows initial improvements but exhibits diminishing returns over time.
GT GEN
Model Metric
No Prior
Single-Day
Multi-Day
No Prior
Single-Day
Multi-Day
Mistral CUI 21.60 ± 6.47 34.06 ± 12.15 30.55 ± 13.48 21.60 ± 6.47 22.30 ± 16.13 22.82 ± 16.47
ROUGE-L 17.70 ± 3.79 26.11 ± 11.68 23.32 ± 13.05 17.70 ± 3.79 17.57 ± 13.47 17.90 ± 13.59
SapBERT 68.16 ± 6.91 71.63 ± 10.86 72.41 ± 9.47 68.16 ± 6.91 70.51 ± 11.60 71.50 ± 9.66
Average 35.82 43.93 42.09 35.82 36.79 37.41
Llama3 CUI 21.83 ± 6.51 45.35 ± 10.45 43.55 ± 12.00 21.83 ± 6.51 40.89 ± 13.61 40.50 ± 13.29
ROUGE-L 16.60 ± 3.39 33.16 ± 10.68 32.42 ± 12.14 16.90 ± 3.85 28.86 ± 12.8 29.34 ± 13.30
SapBERT 72.01 ± 6.04 80.10 ± 8.93 80.42 ± 8.81 72.33 ± 5.48 79.68 ± 10.07 80.48 ± 9.56
Average 36.81 52.87 52.13 37.05 49.81 50.11
Qwen CUI 21.84 ± 5.68 34.27 ± 7.18 34.32 ± 6.65 21.84 ± 5.68 30.10 ± 6.70 29.55 ± 6.50
ROUGE-L 15.35 ± 3.10 77.50 ± 5.37 21.85 ± 4.99 15.18 ± 3.00 18.21 ± 4.28 17.55 ± 4.07
SapBERT 72.57 ± 4.92 20.69 ± 4.37 77.88 ± 4.86 72.97 ± 4.83 78.39 ± 4.11 77.61 ± 4.52
Average 36.87 44.15 44.68 37.12 42.23 41.57
DeepSeek CUI 22.58 ± 6.63 34.45 ± 9.26 34.15 ± 10.00 22.58 ± 6.63 27.33 ± 7.69 27.93 ± 7.46
ROUGE-L 16.16 ± 3.15 21.84 ± 5.65 21.07 ± 5.42 16.16 ± 3.15 17.37 ± 3.86 16.88 ± 3.88
SapBERT 74.97 ± 3.66 78.29 ± 4.61 78.35 ± 5.12 74.97 ± 3.66 77.90 ± 4.36 77.93 ± 4.79
Average 37.90 44.86 44.52 37.90 40.87 40.91
Table 14: Comparison of ground-truth (GT, same results we report in main text) and generated (GEN) settings on assessment and plan generation (results of direct generation approach shown)
Figure 9: Comparison of generated patient notes across different methods of incorporating prior-day context. Each subplot represents a different patient admission, with bars indicating the composition of generated notes across Days 2, 3, and 4. The stacked bars show the proportion of common elements (overlapping between ground truth and generated notes), GT elements (present in the ground truth but missing in generated notes), and GEN elements (newly generated content not found in the ground truth). Method -1 (no prior-day context), Method 1 (single prior-day context), and Method 2 (multi-day context) demonstrate how historical context influences the balance between accurate retention and newly introduced information.
Figure 10: Metric breakdown across modalities for the 24 hour time window (on direct generation)
A.6 More results for using ground-truth progress notes vs. generated progress notes on A&P Generation
This section presents further results comparing ground-truth progress notes and model-generated progress notes as input for A&P Generation. Our analysis evaluates how using prior ground-truth notes versus LLM-generated notes impacts the overall quality, coherence, and accuracy of the generated A&P sections.
Table 14 presents results on all LLMs direct generation. Overall, there are performance decreases when moving from using generated progress notes as context input. The results highlight key differences in information retention and propagation, where models using ground-truth progress notes tend to maintain better clinical consistency, while those using generated notes may accumulate errors over multiple days, leading to drift and hallucination in longitudinal patient summaries. These findings emphasize the need for robust calibration and error correction mechanisms when relying on LLM-generated progress notes for iterative summarization.
Additionally, we include Figure 9 to show the impact of incorporating prior-day context on the composition of generated patient notes. The figure illustrates how different methods—ranging from no prior-day context (Method -1) to single-day (Method 1) and multi-day context (Method 2)—affect the alignment between generated elements (GEN), ground truth elements (GT), and common elements across multiple admissions. As the number of prior days included increases, we observe changes in the proportion of correctly retained ground truth elements and newly generated content, highlighting the role of historical context in improving consistency and completeness in patient note generation. Additionally, the inclusion of prior context reduces the number of newly generated elements, indicating a shift away from generating potentially extraneous or less relevant content.
However, the degree of improvement varies across different admissions, suggesting that some cases benefit more from historical context than others, potentially due to differences in case complexity or the structure of prior notes. Furthermore, the differences between generated and ground truth elements become more pronounced as days progress, highlighting the challenge of maintaining consistency in patient notes over time. Overall, the findings suggest that incorporating multi-day context enhances the accuracy and stability of generated patient notes, reducing hallucinated content while preserving clinically relevant information.
Physician Evaluation of LLM-Generated Clinical Summaries
Thank you for participating in this evaluation. Please assess the LLM-generated text based on the following criteria using the 5-point Likert scale provided for each question.
Scale Definitions (from 1-5):
1 Strongly Disagree; 2 Disagree; 3 Neutral; 4 Agree; 5 Strongly Agree
Category Evaluation Question Score (1-5)
Overall Accuracy Xu et al. (2024a); Singhal et al. (2023); Ben Abacha et al. (2023); Croxford et al. (2024); Johri et al. (2025) How well does the generated text align with the actual clinical data?
The summary is factually correct and accurately represents the original data.
No major distortions or misinterpretations of key clinical facts.
Hallucination (Faithfulness) Ben Abacha et al. (2023); Singhal et al. (2023); Aljamaan et al. (2024); Johri et al. (2025) To what extent does the LLM generate information faithfully?
The generated text does not introduce any fabricated, misleading, or incorrect information.
All statements in the summary can be traced back to the original document.
Omission Croxford et al. (2024); Ben Abacha et al. (2023) Did the model include all important clinical details?
The generated summary includes all clinically important details.
No missing critical pieces of information relevant to patient care.
Readability Xu et al. (2024a) How easy is the generated text to read and comprehend?
The text is well-structured, clear, and easy to understand.
Uses appropriate medical terminology without excessive complexity.
Clinical Relevance (Usefulness) Singhal et al. (2023); Johri et al. (2025) How useful is the content for clinical decision-making?
The information provided is highly relevant to the clinical task.
Avoids unnecessary or unrelated details.
Specificity (Level of Detail) Williams et al. (2024) Does the summary maintain an appropriate level of detail?
The text balances between a high-level summary and necessary details.
Avoids being overly vague or excessively detailed.
Table 15: Physician evaluation survey
A.7 Break-down results behind Figure 4
Recall that Figure 4 aggregates all metrics for LLM direct generation on discharge summarization. To provide comprehensive results analysis, we include Figure 10 for 24h window and 48h window. The trends of performance changing across different modalities are consistent with the Figure 4.
A.8 Details about human evaluation
Last but not least, we present the full survey questions aggregated from existing work in Table 15. We would like to emphasize that this survey has not been validated, rather, it selects the criteria after consulting with a physician regarding what they value in LLM generated text and findings from prior work, as cited in the table. These questions allow us to evaluate the accuracy, faithfulness, readability, and clinical relevance of LLM-generated text, providing a structured framework for assessing their strengths and limitations in a clinical setting.